OpenStudy (saldis):

Solve for x. 7x=4x+277x=4x+27 Enter your answer in the box. x =

OpenStudy (f_jayyy):

its not possible to have two equal signs in one equation

OpenStudy (saldis):

lol it repeated its self

OpenStudy (f_jayyy):

so what is the right equation so i can solve it

OpenStudy (f_jayyy):

i will be bck in 10 mins

OpenStudy (saldis):

Solve for x. 7x=4x+27 Enter your answer in the box. x =

OpenStudy (f_jayyy):

7x - 4x = 27 3x = 27 3x/3 = 27/3 x = 9